### v1

@@ C.3 Performance Measures
where $q$ denotes the forecast horizon, $\widehat{Y}(x,{T+v})$ is the $v^{th}$-step ahead forecast based on $T$ historical data for the $x^{th}$ prefecture, and $Y(x,{T+v})$ represents the corresponding ground truth.
C.4 Statistical Significance
@@
To validate the robustness of the proposed EGDL-Parallel and EGDL-Series frameworks, we employ the multiple comparisons with the best (MCB) test [56]. This model-agnostic procedure identifies the framework with the lowest average rank as the ‘best’ performing model. To compare the relative performance of the architectures, the MCB test calculates a critical distance (CD) based on the Tukey distribution and utilizes the CD of the ‘best’ model as the reference threshold. The MCB test results, based on the SMAPE metric, comparing EGDL-Parallel and EGDL-Series forecasters with baseline models, are visualized in Fig. A6. Among the EGDL-Parallel architectures, the EGP-NHits framework achieves the lowest rank of 1.00, followed by the NHits and EGP-NBeats frameworks. Similarly, for the EGDL-Series architectures, the EGS-NHits model has the lowest rank of 2.00, followed by EGS-TCN and NHits. These results indicate that the epidemic-guided NHits models consistently deliver the ‘best’ performance among the evaluated frameworks. The CD for the ‘best’ models, highlighted by the shaded regions, serves as the reference value in the MCB test. The competing architectures with CDs exceeding this reference value are deemed to perform significantly worse than the epidemic-guided NHits frameworks. Notably, all epidemic-guided forecasting approaches exhibit lower average ranks compared to their respective baseline architectures, underscoring the effectiveness of incorporating epidemic principles into data-driven forecasting models.
@@
Figure A6: The MCB test results, comparing (a) EGDL-Parallel approaches and (b) EGDL-Series architectures with baseline models using the SMAPE metric. In the plots, the Y-axis represents the average ranks of the models, while the X-axis labels each model along with its corresponding average rank. For instance, the label ‘EGP-NHits - 1.00’ indicates that the EGP-NHits model achieved an average rank of 1.00, similar for other approaches.
C.5 Uncertainty Quantification in EGDL Approaches
Alongside the point forecasts generated by the EGDL frameworks, we quantify the uncertainties associated with these predictions using the conformal prediction approach [60]. This distribution-free method constructs prediction intervals at a pre-specified confidence level, providing robust uncertainty estimates. The prediction intervals of the EGDL-based forecasters make it a suitable choice for real-world decision-making. The conformal prediction process begins by fitting an uncertainty model to the input data and calculating conformal scores, which are derived from the residuals of the EGDL frameworks scaled by the uncertainty model’s predictions. Using these scores, conformal quantiles are computed via a weighted aggregation technique. The prediction intervals are then determined by adding (subtracting) the EGDL forecasts with the upper (lower) quantiles scaled by the uncertainty model predictions. Fig. A7 presents the conformal prediction intervals for the EGS-TCN model across selected prefectures of Japan, alongside the forecasts of baseline models. As the plots highlight, the prediction intervals effectively capture the TB incidence trends in most scenarios, demonstrating superior generalization compared to the baseline and networked SIR models. These probabilistic intervals demonstrate the forecast’s uncertainties and are crucial for designing effective disease intervention strategies.
Figure A7: Visualization of the ground truth (red dots) tuberculosis (TB) incidences monitored at prefectures 01 (left) and 27 (right) from Jan - Dec 2015 with the corresponding point forecasts of EGS-TCN (blue line), TCN (green line), the infected curve of the networked SIR framework (violet line), and conformal prediction interval (yellow-shaded) of EGS-TCN.
Data and Code Availability Statement

### v2

@@ C.2 Performance Measures
where $q$ denotes the forecast horizon, $\widehat{Y}(x,{T+v})$ is the $v^{th}$-step ahead forecast based on $T$ historical data for location $x$, and $Y(x,{T+v})$ represents the corresponding ground truth.
C.3 Statistical Significance
@@
To validate the robustness of the proposed EGDL-Parallel and EGDL-Series frameworks, we employ the multiple comparisons with the best (MCB) test [61]. This model-agnostic procedure identifies the framework with the lowest average rank as the ‘best’ performing model. To compare the relative performance of the architectures, the MCB test calculates a critical distance (CD) based on the Tukey distribution and utilizes the CD of the ‘best’ model as the reference threshold. The MCB test results, based on the SMAPE metric, comparing EGDL-Parallel and EGDL-Series forecasters with baseline models, are visualized in Fig. A5 and A6 for Japan and China, respectively. For Japan’s TB dataset, among the EGDL-Parallel architectures, the EGP-NHits framework achieves the lowest rank of 1.00, followed by the NHits and EGP-NBeats frameworks. Similarly, for the EGDL-Series architectures, the EGS-NHits model has the lowest rank of 2.00, followed by EGS-TCN and NHits. In case of TB incidence cases of China, the EGP-NHits and EGS-NBeats frameworks achieve the minimum ranks of 2.50 and 2.00, respectively. These results indicate that the epidemic-guided NHits and NBeats models consistently deliver the ‘best’ performance among the evaluated frameworks. The CD for the ‘best’ models, highlighted by the shaded regions, serves as the reference value in the MCB test. The competing architectures with CDs exceeding this reference value are deemed to perform significantly worse than the ‘best’ performing frameworks. Notably, all epidemic-guided forecasting approaches exhibit lower average ranks compared to their respective baseline architectures, underscoring the effectiveness of incorporating epidemic principles into data-driven forecasting models.
@@
Figure A5: The MCB test results, comparing (a) EGDL-Parallel and (b) EGDL-Series architectures with baseline models using the SMAPE metric. In the plots, the Y-axis represents the average ranks of the models, while the X-axis labels each model along with its corresponding average rank. For instance, the label ‘EGP-NHits - 1.00’ indicates that the EGP-NHits model achieved an average rank of 1.00, similar to other approaches. A lower rank indicates a better-performing model based on the evaluation metric used.
Figure A6: The MCB test results, comparing (A) EGDL-Parallel and (B) EGDL-Series architectures with baseline models using the SMAPE metric for forecasting TB incidence cases of China. In the plots, the Y-axis represents the average ranks of the models, while the X-axis labels each model along with its corresponding average rank. For instance, the label ‘EGP-NHits - 2.50’ indicates that the EGP-NHits model achieved an average rank of 2.50, similar to other approaches. A lower rank indicates a better-performing model based on the evaluation metric used.